OpenStudy (anonymous):

Use the three steps to solve the problem. The sum of 3 consecutive integral numbers is 117. Find the numbers.

zepdrix (zepdrix):

"Use the three steps to solve the problem." What 3 steps do they want you to use? c:

OpenStudy (anonymous):

Step 1: Let x = small number. Then from the first sentence, 4x = large number. Step 1 is complete as the two quantities to be found are assigned variables. Step 2: The second sentence states conditions placed on the quantities: "their sum is 45." So, x + 4x = 45. Step 3: Solve x + 4x = 45. 5x = 45 x = 9(small number) 4x = 36(large number) The solution set to this number problem is {9, 36}.

zepdrix (zepdrix):

Oh, so algebraically :) with x n stuff. Hmm ok.

OpenStudy (anonymous):

ya lol

zepdrix (zepdrix):

So if we let the smallest of the 3 numbers be X, SINCE they are consecutive (one number right after the next), We can write the number AFTER X as (X+1) and the number after that as (X+2) Does that part make sense? :)

OpenStudy (anonymous):

kinda :P

zepdrix (zepdrix):

if x = 20 20+1= x + 1 Which is the next number in the sequence :D we're doing that type of thing.

zepdrix (zepdrix):

uh oh kaite still confused? c: These x's messin with yer head? :D

OpenStudy (anonymous):

:D it sure is i hate math

zepdrix (zepdrix):

aw :c maf is awesome

OpenStudy (anonymous):

not to me :(

zepdrix (zepdrix):

Step 2, hmm so we recognize the conditions placed on our quantities. "the sum is 117". So if X= a number. (let's say ummm 37 how bout.) Then we can write 38 (the next number after 37) 37+1 = x+1 And we can write 39 (the number after 38) 37+2 = x+2 So the sum of our 3 numbers will be, 37+38+39=117 37 + (37+1) + (37+2) = 117 x + (x+1) + (x+2) = 117 Add up the x's 3x + 3 = 117 Then solve for x! Normally you wouldn't know that 37 38 and 39 are your numbers ahead of time. But I'm trying to keep it simple so you can maybe grasp this concept <:o

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

actually that only equals 114 so would't it be 38 39 40

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Oh appearantly it was 38, 39 and 40! ah sorry bout that! XD

OpenStudy (anonymous):

it okay